# AI-Generated Images Introduce Invisible Relevance Bias to Text-Image Retrieval

### 23 Nov 2023 v1

@@ Abstract.
With the advancement of generation models, AI-generated content (AIGC) is becoming more realistic and less discernible, flooding the Internet. A recent study suggests that this phenomenon has elevated the issue of source bias in text retrieval for web searches. Specifically, neural retrieval models tend to rank generated texts higher than human-written texts.
@@ 1. Introduction
Figure 1. Bias found in our paper. IR models tend to rank AI-generated images higher than real images even though they have very similar semantics. This bias increases the likelihood of the generated images being extensively exposed within vast datasets, which makes them more likely to be mixed into the training of AIGC and retrieval models, leading to more serious bias and forming a vicious cycle.
@@
Information retrieval (IR) is an important way for people to obtain the target information from massive Internet information (Manning, 2009). However, the rapid proliferation of AI-generated content (AIGC) presents a significant new challenge to IR. Specifically, as the Internet becomes increasingly populated with AI-generated content, the question arises: How will these contents influence the ranking results of retrieval systems? In response to this challenge, (Dai et al., 2023) conducted a study on text modality and found that neural information retrieval models exhibit a preference for texts generated by Large Language Models (LLMs). They refer to this category of biases in neural retrieval models towards the LLM-generated text as the source bias. However, beyond textual information, the Internet is replete with a substantial number of images, serving as crucial sources for IR systems. Building on this, our paper extends the investigation of source bias in AI-generated content to text-image retrieval models. A pivotal question emerges: What impact will AI-generated images have on existing text-image retrieval models?
@@
A crucial distinction from research on text modality lies in the methodology for constructing a benchmark to simulate retrieval scenarios comprising both real images and AI-generated images for IR models. A reasonable scenario for assessing the potential bias requires that the generated images and the real images have sufficiently similar visual semantics. This can avoid increasing or decreasing some semantic associations between the generated images and the query caused by the introduction or lack of some semantic information during the image generation. In other words, the IR model preferring (or rejecting) an AI-generated image that is more (or less) semantically similar to the query than the real image cannot conclusively prove the existence (or nonexistence) of the bias. Study on text modality prompts LLMs to rewrite each real text in the corpus, thereby constructing an LLM-generated text for each real text without introducing additional semantic information. However, this approach is unsuitable for constructing the data in image modality. It is because we find mainstream image generation models such as diffusion models (Rombach et al., 2022) are not good at following prompts to rewrite images while maintaining semantics. Besides, the rewriting paradigm diverges from the prevalent content generation scenario, where receiving a description as a prompt and generating the texts or images. To solve the above problem, we propose an image over-sampling and selection strategy based on the generation according to the merged caption. Experimental results and human evaluation show that our proposed method can successfully construct semantically similar AI-generated images for real images. The mixture of these images provides an effective simulation and validation benchmark for investigating text-image retrieval models within scenarios featuring a mix of AI-generated and real images. Our experiments on this benchmark reveal that text-image retrieval models tend to rank the AI-generated images higher than the real images, even though the AI-generated images do not exhibit more visually relevant features to the query than real images. We define this bias as invisible relevance bias introduced by AI-generated images.
Another important point that is not considered in previous work is that AI-generated content does not only have an impact on the inference stage of the retrieval models. Due to their large-scale distribution on the Internet, they are very likely to be mixed into the training data of retrieval models in the future. Therefore, we further mix the AI-generated images into the training of the retrieval model and explore the rules of its impact on the retrieval models by adjusting the mixing ratio. Our experimental results show that as the mixing ratio of AI-generated images increases, the invisible relevance bias becomes more serious, and the retrieval performance of on the generated images gradually becomes stronger, while the retrieval performance on the real images is gradually damaged. This phenomenon reveals that invisible relevance bias causes the generated images to have a higher chance of being obtained from massive data, which makes them more likely to be mixed into the training of AIGC and retrieval models (Shumailov et al., 2023), leading to more serious bias and forming a vicious cycle. (shown in Figure 1)
To mitigate the bias, we propose an effective training method for debiasing. Specifically, we introduce a contrastive loss between real and AI-generated images during training. This contrastive loss measures the invisible relevance bias introduced by the AI-generated images for retrieval models. Optimization of this loss can reduce the invisible relevance bias, thereby alleviating the preference for AI-generated images in text-image retrieval models. In addition, we also introduce the sampling probability to enable our debiasing method to dynamically adjust the preference to AI-generated images of text-image retrieval models.
Last but not least, we use our proposed debiasing method to reversely determine that the causes of invisible relevance is the generated images causing the image encoder to embed additional information into their representation. This information has a certain consistency in different generated images and can make the retriever estimate a higher relevance score. We also design experiments to support this.
The contributions of this paper can be concluded as:
(1) We reveal the potential impact of AI-generated images on the ranking results of text-image retrieval systems in the context of the rapid development of AIGC. To reasonably study the impact, we propose an effective method to construct a reasonable benchmark to simulate text-image retrieval in scenarios consisting of real and AI-generated images.
@@
(2) Experimental results show that AI-generated images introduce the invisible relevance to text-image retrieval models, causing the bias that retrieval models prefer ranking AI-generated images higher even though the AI-generated images do not exhibit more visually relevant features to the query than real images. This bias widely exists in retrieval models with different training data and architectures, from models trained from scratch to models pre-trained on massive image-text pairs, from dual-encoder retrieval models to fusion-encoder retrieval models. We also reveal that the loop formed by training and inference causes the retrieval model to fall into a vicious cycle of increasingly serious bias.
@@ 2. Benchmark Construction
This paper focuses on revealing the potential impact of AI-generated images on the ranking results of text-image retrieval systems. The prerequisite for all this is to construct a retrieval database consisting of real images and AI-generated images. This database should meet four critical requirements. (-$\mathcal{R}\mbox{-}1$) Firstly, from the perspective of fair assessment, a reasonable scenario for assessing the potential bias requires that the generated images and the real images must have sufficiently similar visual semantics. This can avoid increasing or decreasing some semantic associations between the generated images and the query caused by the introduction or lack of some semantic information during the image generation. That is, the IR model preferring (or rejecting) an AI-generated image that is more (or less) semantically similar to the query than the real image cannot prove the existence (or nonexistence) of the bias. (-$\mathcal{R}\mbox{-}2$) Secondly, retrieval performance on the database containing only generated images should not change significantly compared to retrieval performance on real images only. This can further ensure that no additional visual semantics relevant (or irrelevant) to the query are introduced during the image generation. (-$\mathcal{R}\mbox{-}3$) Thirdly, the image generation pipeline should be in line with the most common scenario of content generation that receives a description as prompt and generates the texts or images, so that the obtained bias assessment results can be more consistent with the actual scene. (-$\mathcal{R}\mbox{-}4$) Fourthly, the number of generated images and real images should be equal, so as to avoid additional bias caused by the difference in the number of images. Based on the above three requirements, we propose an image over-sampling strategy based on the generation via merged caption, which will be introduced in the following contents.
2.1. Image Generation
2.1.1. Generation Based on Single Caption
In this method, we randomly select one caption from the five captions of each real image and use the selected caption as the prompt of the stable diffusion model [^1] https://huggingface.co/stabilityai/stable-diffusion-xl-base-1.0 to obtain the AI-generated image corresponding to the real image.
2.1.2. Generation Based on Merged Caption
Since a single caption is not guaranteed to contain the complete visual semantics of the image, we propose to merge five captions to obtain a relatively complete semantic description of the image. Specifically, for each real image ($I_{r}$), we use gpt-3.5-turbo to combine the five captions to get a new merged caption. And input the merged caption to a stable diffusion model $\mathbf{M}$ to generate the image ($I_{g}$), which can be described as:
[…]
$I^{g}$ $=\mathbf{M}(C_{m}).$
The intention of this design is that different captions of an image may be the descriptions from different perspectives of the image. Therefore, merging these captions can obtain an overall description of the image, thereby making stable diffusion generate images that have sufficiently similar visual semantics to the real image as much as possible. Besides, this also avoids the relevance bias in assessment caused by the generated image being overly inclined to one certain caption.
2.1.3. Generation Based on Image Rewriting
We not only input the merged caption $C_{m}$ but also input the real image $I_{r}$ as the initial image to condition the generation of the new image. Generation constrained by the input real image can output the AI-generated image that has higher similarity with the real image.
@@ 2.1.3. Generation Based on Image Rewriting
Table 1. Comparison between different generation methods under four requirements mentioned at the beginning of Section 2. Similarity is the cosine between embeddings of real and generated images encoded by OpenClip. Retrieval Performance is the difference in retrieval performance of BEIT-3 on the databases only containing generated images and only containing real images respectively.
@@
2.1.4. Generation Based on Image Over Sampling and Selection
In order to further narrow the semantic similarity between the generated images and real images, we propose an image over-sampling strategy. Specifically, for a real image $I^{r}$, we use stable diffusion to perform multiple times generation with different random seeds and get $n$ generated images $\{{I^{g}_{1}},{I^{g}_{2}},...,{I^{g}_{n}}\}$. Then, we use the vision encoder $v(\cdot;\theta)$ of a powerful open source pre-trained vision-language model [^2] https://huggingface.co/laion/CLIP-ViT-H-14-laion2B-s32B-b79K to get the embedding $\bm{e_{r}}$ for $I^{r}$ and set of embeddings $\bm{E}=\{\bm{{e^{g}_{1}}},\bm{{e^{g}_{n}}},...,\bm{{e^{g}_{n}}}\}$ for $\{{I^{g}_{1}},{I^{g}_{2}},...,{I^{g}_{n}}\}$. We calculate the cosine similarity between each embedding in $\bm{E}$ and $\bm{e_{r}}$ to get the similarity between the generated images and the real image $I^{r}$. Finally, we select the generated image with the maximum similarity between $I^{r}$ as the final generated image ${I}^{g}$ for the real image:
(1) $I^{g}=\arg\max_{i}\{dot(\bm{e^{r}},\bm{e^{g}_{i}})|\bm{{e^{g}_{i}}}\in\bm{E}\}.$
2.1.5. Comparison between Different Generation Methods
The comparison between different generation methods under the above-mentioned four requirements is shown in Table 1. Taking these results into account, we choose generation based on merged caption and image over-sampling and selection as the image generation method (shown in Figure 2) to construct our benchmark. Using this method, we can generate the images for each real image in Flicker30k and MSCOCO. For the convenience of description, in the following contents, Flicker30k and MSCOCO indicate the datasets that only contain real images, Flicker30k (AIGC) and MSCOCO (AIGC) indicate the datasets that only contain AI-generated images, Flicker30k+AIGC and MSCOCO+AIGC indicate the datasets that consist of both real and AI-generated images, each real image has an AI-generated image.
Figure 2. Pipeline of our benchmark construction.
@@ 2.2. Assessing the Quality of Generated Images
Generated images and real images having sufficiently similar visual semantics so as to avoid introducing more semantics relevant to the query is a prerequisite for fairly assessing the bias caused by invisible relevance. In this section, we propose two methods to measure the semantic similarity. Firstly, we perform retrieval on two databases that only contain real images and only contain generated images respectively, and count the retrieval performance. The intention for this is that if the generated images have more semantics relevant to their corresponding queries than the real images, images in the database will become easier to distinguish, resulting in the retrieval performance on the database that only contains generated images significantly higher than the performance on the database that only contains real images. Secondly, we introduce human evaluation to further evaluate the semantic information from human vision. Specifically, we let humans determine whether the generated images have more visual semantics relevant to the queries compared to the real images.
@@ 2.2.1. Retrieval Performance
Table 2. Performance of the retrieval models on the databases that only contain real images and only include AI-generated images respectively. Performance on the AI-generated images is not significantly better than the performance on real images can demonstrate the success of our benchmark construction strategy.
Flicker30k or Flicker30k (AIGC) MSCOCO or MSCOCO (AIGC)
NDCG@1 NDCG@3 NDCG@5 R@1 R@3 R@5 NDCG@1 NDCG@3 NDCG@5 R@1 R@3 R@5
FLAVA Real 38.70 45.72 47.26 63.12 82.52 88.53 36.65 44.45 46.27 59.77 81.38 88.62
AI-generated 41.59 47.45 48.72 67.83 84.10 89.10 36.49 43.29 45.17 59.52 78.44 85.90
ALIGIN Real 45.43 50.62 51.65 74.10 88.38 92.44 38.13 44.71 46.33 62.20 80.52 87.00
AI-generated 43.47 49.38 50.42 70.90 87.24 91.38 36.34 43.32 45.06 59.28 78.62 85.51
BEIT-3 Real 47.45 52.15 52.87 77.40 90.32 93.20 41.24 47.16 48.63 67.25 83.72 89.52
AI-generated 45.31 50.49 51.43 73.90 88.24 91.94 38.33 44.76 46.19 62.52 80.30 85.98
We introduce three open-source widely used and powerful vision-language pre-trained models including FLAVA (Singh et al., 2022), ALIGN (Jia et al., 2021), and BEIT-3 (Wang et al., 2022) to perform retrieval on the databases that only contain real images and only contain generated images respectively. Since text-image retrieval datasets such as Flicker30k and MSCOCO are composed of real images, the training on these images will further introduce additional bias. In view of the fact that the above three models have shown excellent performance in zero-shot text-image retrieval (Wang et al., 2022), we directly use the above three models to perform retrieval under the zero-shot setting. The experimental results in Table 2 show that retrieval performance on the database containing only AI-generated images is not significantly greater than retrieval performance on the database containing only real images. It indicates that the AI-generated images in our benchmark do not introduce more visual semantics relevant to the query.

### 15 Jan 2024 v3

@@ Abstract.
With the advancement of generation models, AI-generated content (AIGC) is becoming more realistic and less discernible, flooding the Internet. A recent study suggests that this phenomenon causes source bias in text retrieval for web search. Specifically, neural retrieval models tend to rank generated texts higher than human-written texts.
@@ 1. Introduction
Information retrieval (IR) is an important way for people to obtain the target information from massive Internet information (Manning, 2009). However, the rapid proliferation of AI-generated content (AIGC) presents a significant new challenge to IR: as the Internet becomes increasingly inundated with AI-generated contents, how will these contents influence the ranking results of search? In response to this challenge, (Dai et al., 2023) conducted a study on text modality and found that neural information retrieval models exhibit a preference for texts generated by Large Language Models. They refer to this category of biases in neural retrieval models towards the LLM-generated text as the source bias. However, beyond textual information, the Internet is replete with a substantial number of images, serving as crucial sources for IR systems. Building on this, our paper extends the investigation of source bias in AI-generated content to text-image retrieval models. A pivotal question emerges: What impact will AI-generated images have on existing text-image retrieval models?
Figure 1. Bias found in our paper. IR models tend to rank AI-generated images higher than real images even though they have very similar semantics. This bias increases the likelihood of the generated images being exposed from massive data of Internet, which makes them more likely to be mixed into the training of AIGC and retrieval models, leading to more serious bias and forming a vicious cycle.
@@
A crucial distinction from research on text modality lies in the methodology for constructing a benchmark to simulate retrieval scenarios comprising both AI-generated and real images for IR models (§ 2). A reasonable scenario for assessing the potential bias requires that the generated images and the real images have sufficiently similar visual semantics. This can avoid increasing or decreasing some semantic associations between the generated images and the query caused by the image generation. In other words, the IR model preferring (or rejecting) an AI-generated image that is more (or less) semantically relevant to the query than the real image cannot conclusively prove the existence (or nonexistence) of the bias. Study on text modality prompts LLMs to rewrite each real text in the corpus, thereby constructing an LLM-generated text for each real text without introducing additional semantic information. However, this approach is unsuitable for constructing the data in image modality. It is because we find mainstream image generation models such as diffusion models (Rombach et al., 2022) are not good at following prompts to rewrite images while maintaining semantics. Besides, the rewriting paradigm diverges from the prevalent content generation scenario that receives a description as a prompt and generates the texts or images. To solve the above problem, we propose an image over-sampling and selection strategy based on the merged caption (§ 2.1). Experimental results and human evaluation show that our proposed method can successfully construct semantically similar AI-generated images for real images (§ 2.2). The mixture of these images provides an effective simulation benchmark for investigating text-image retrieval models within scenarios featuring a mix of AI-generated and real images.
@@
Our experiments on the constructed benchmark reveal that text-image retrieval models tend to rank the AI-generated images higher than the real images, even though the AI-generated images do not exhibit more visually relevant semantics to the queries than real images (§ 3.3). We define this bias as invisible relevance bias introduced by AI-generated images. This bias widely exists in retrieval models with different training data and architectures, including models trained from scratch and models pre-trained on massive image-text pairs, and encompassing dual-encoder and fusion-encoder retrieval models.
Another important point that is not considered in previous work is that AI-generated content does not only have an impact on the inference stage of the retrieval models. Due to their wide distribution on the Internet, they are very likely to be mixed into the training data of retrieval models in the future. Therefore, we further mix the AI-generated images into the training data of the retrieval model and explore the rules of its impact on the retrieval models by adjusting the mixing ratio. Our experiments show that as the mixing ratio of AI-generated images in training data increases, the invisible relevance bias becomes more serious (§ 3.4), the retrieval model exhibits a greater inclination to rank generated images higher. This phenomenon reveals that invisible relevance bias causes the generated images to have a higher chance of being obtained from massive data, which makes them more likely to be mixed into the training of generation and retrieval models and causes the model collapse (Shumailov et al., 2023), leading to more serious bias and forming a vicious cycle. (shown in Figure 1)
To mitigate the bias, we propose an effective training method for debiasing (§ 4). Specifically, we introduce a contrastive loss between real and AI-generated images during training. This contrastive loss measures the invisible relevance bias introduced by the AI-generated images for retrieval models. Optimization of this loss can reduce the invisible relevance bias, thereby alleviating the preference to AI-generated images in text-image retrieval models. In addition, we also introduce the sampling probability to enable our debiasing method to dynamically adjust the tolerance to AI-generated images of text-image retrieval models.
Last but not least, we use our proposed debiasing method to reversely determine that the causes of invisible relevance bias is the generated images causing the image encoder to embed additional information into their representation (§ 5). This additional information has a certain consistency in different generated images and can make the retriever estimate a higher relevance score. We also design experiments to support this.
The main contributions of this paper can be concluded as:
(1) We reveal the potential impact of AI-generated images on the ranking results of text-image retrieval systems in the context of the rapid development of AIGC. To reasonably study the impact, we construct a reasonable benchmark to simulate text-image retrieval in scenarios consisting of both real and AI-generated images.
(2) Experimental results show that AI-generated images introduce the invisible relevance to text-image retrieval models, causing the bias that retrieval models prefer ranking AI-generated images higher even though the AI-generated images do not exhibit more visually relevant semantics to the queries than real images. This bias widely exists in retrieval models with different training data and architectures. We also reveal that the loop formed by training and inference causes the retrieval model to fall into a vicious cycle of increasingly serious bias.
@@ 2. Benchmark Construction
This paper focuses on revealing the potential impact of AI-generated images on the ranking results of text-image retrieval systems. The prerequisite for all this is to construct a retrieval database consisting of real images and AI-generated images. This database should meet four critical requirements. (-$\mathcal{R}\mbox{-}1$) Firstly, from the perspective of fair assessment, a reasonable scenario for assessing the potential bias requires that the generated images and the real images must have sufficiently similar visual semantics. This can avoid increasing or decreasing some semantic associations between the generated images and the query caused by image generation. That is, the IR model preferring (or rejecting) an AI-generated image that is more (or less) semantically relevant to the query than the real image cannot prove the existence (or nonexistence) of the bias. (-$\mathcal{R}\mbox{-}2$) Secondly, retrieval performance on the database containing only generated images should not change significantly compared to retrieval performance on real images only. This can further ensure that no additional visual semantics relevant (or irrelevant) to the query are introduced during the image generation. (-$\mathcal{R}\mbox{-}3$) Thirdly, the image generation pipeline should be in line with the most common scenario of content generation that receives a description as prompt and generates the texts or images, so that the obtained bias assessment results can be more consistent with the actual scene. (-$\mathcal{R}\mbox{-}4$) Fourthly, the number of generated images and real images should be equal to achieve fair comparison.
@@
Table 1. Compare generation methods based on the four requirements outlined in Section 2. Similarity is the cosine between embeddings of real and generated images encoded by OpenClip. Retrieval Performance is the difference in retrieval performance of BEIT-3 on the databases only containing generated images and only containing real images respectively.
@@
2.1. Image Generation
Generation Based on Single Caption. In this method, we randomly select one caption from the five captions of each real image and use the selected caption as the prompt of stable diffusion model [^1] https://huggingface.co/stabilityai/stable-diffusion-xl-base-1.0 to get the AI-generated image corresponding to the real image.
Generation Based on Merged Caption. Since a single caption may not contain the complete visual semantics of the image, we propose to merge five captions to obtain a relatively complete semantic description of the image. Specifically, for each real image ($I_{r}$), we use gpt-3.5-turbo to combine the five captions to get a newly merged caption. And input the merged caption to a stable diffusion model $\mathbf{M}$ to generate the image ($I_{g}$), which can be described as:
$C_{m}=\textrm{GPT-3.5}(C_{1},C_{2},...,C_{n}),\quad I^{g}=\mathbf{M}(C_{m}).$
The intention of this design is that different captions of an image can be the descriptions from different perspectives of the image. Therefore, merging these captions can obtain an overall description of the image, thereby making stable diffusion generate images that have sufficiently similar visual semantics to the real image. Besides, this also avoids the relevance bias in assessment caused by the generated image being overly inclined to one certain caption.
Generation Based on Image Rewriting. In addition to the merged caption $C_{m}$, the real image $I_{r}$ is also used as the initial image to condition the generation of the new image. Generation constrained by the input real image can output the generated image that has higher similarity to the real image. But it also puts forward higher requirements for the multi-modal understanding ability of the model.
Generation Based on Image Over Sampling and Selection. To further narrow the semantic similarity between the generated images and real images, we propose an image over-sampling strategy. Specifically, for a real image $I^{r}$, we use stable diffusion to perform multiple times generation with different random seeds and get $n$ generated images $\{{I^{g}_{1}},{I^{g}_{2}},...,{I^{g}_{n}}\}$. Then, we use the vision encoder $v(\cdot;\theta)$ of a powerful open source pre-trained vision-language model [^2] https://huggingface.co/laion/CLIP-ViT-H-14-laion2B-s32B-b79K to get the embedding $\bm{e_{r}}$ for $I^{r}$ and set of embeddings $\bm{E}=\{\bm{{e^{g}_{1}}},\bm{{e^{g}_{2}}},...,\bm{{e^{g}_{n}}}\}$ for $\{{I^{g}_{1}},{I^{g}_{2}},...,{I^{g}_{n}}\}$. We calculate the cosine similarity between each embedding in $\bm{E}$ and $\bm{e_{r}}$ to get the similarity between the generated images and the real image $I^{r}$. Finally, we select the generated image with the maximum similarity between $I^{r}$ as the final generated image ${I}^{g}$ for the real image:
(1) $I^{g}=\arg\max_{i}\{cosine(\bm{e^{r}},\bm{e^{g}_{i}})|\bm{{e^{g}_{i}}}\in\bm{E}\}.$
Comparison between Different Generation Methods. The comparison between different generation methods under the above-mentioned four requirements is shown in Table 1. Taking these results into account, we choose generation based on merged caption and image over-sampling and selection as the image generation method to construct our benchmark. Using this method, we can get an AI-generated image for each real image in Flicker30k and MSCOCO. For the convenience of description, in the following contents, Flicker30k and MSCOCO indicate the datasets that only contain real images, Flicker30k (AI) and MSCOCO (AI) indicate the datasets that only contain AI-generated images, Flicker30k+AI and MSCOCO+AI indicate the datasets that consist of both real and AI-generated images..
@@ 2.2. Assessing the Quality of Generated Images
Ensuring sufficiently similar visual semantics between generated and real images to avoid increasing or decreasing relevance to the query is a prerequisite for assessing the bias. In this section, we propose two methods to evaluate this. Firstly, we perform retrieval on the databases that only contain real images and only contain generated images and count the retrieval performance respectively. The intention for this is that if the generated images have more semantics relevant to their corresponding queries than the real images, distinguishing images becomes easier, resulting in significantly higher retrieval performance on the generated images database than on the real images database. Secondly, we introduce human evaluation to further evaluate the semantic information from human vision. Specifically, we let humans determine whether the generated images have more visual semantics relevant to the queries compared to the real images.
@@ 2.2.1. Retrieval Performance
Table 2. Retrieval performance (NDCG) on the databases that only contain real images and only contain AI-generated images. Performance on AI-generated images is not significantly better than the performance on real images can demonstrate the success of our benchmark construction.
Flicker30k or Flicker30k (AI) MSCOCO or MSCOCO (AI)
N@1 N@3 N@5 N@1 N@3 N@5
FLAVA Real 38.70 45.72 47.26 36.65 44.45 46.27
AI 41.59 47.45 48.72 36.49 43.29 45.17
[…]
AI 43.47 49.38 50.42 36.34 43.32 45.06
[…]
AI 45.31 50.49 51.43 38.33 44.76 46.19
The experimental results in Table 2 show that retrieval performance on the database containing only AI-generated images is not significantly greater than retrieval performance on the database containing only real images. It indicates that the AI-generated images in our benchmark do not introduce more visual semantics relevant to the query. Specifically, we use three open-source and powerful vision-language pre-trained models including FLAVA (Singh et al., 2022), ALIGN (Jia et al., 2021), and BEIT-3 (Wang et al., 2022) to perform retrieval on the databases that only contain real images and only contain generated images respectively. Since text-image retrieval datasets are composed of real images, the training on these images will introduce additional bias. Considering the excellent zero-shot text-image retrieval performance of the three models (Wang et al., 2022), we use these models directly for retrieval in zero-shot setting.